How many milimoles of citric acid H3A must be added to 100mL of 0.200M sodium citrate (Na3A) to give a buffer of pH 7.00? pK1=3.13, pK2=4.76, pK3=6.40
Answer #1

we have tribasic acid H3A the dissociation constant

H3A = H2A + H+ is k1

least the pKa more will be the Ka value as per formula

pKa = - log Ka

we know for poly basic acids the dissociation constant which ever is greatest of all always for the removal of first proton is considered as dissociation constants.= k1.

we know the formula for pH of buffer of a weak acid with its highly dissociative salt

pH = pk1 + log [salt]/ [acid]

mili moles of salt present = 100 x 0.2 = 20 milimoles

let milimoles of acid are = n

7.00 = 3.13 + log [ 20 ] / [mili moles of acid ]

log [20/n] = 3.87

20 / n = 103.87 = 7413.1

n = mili moles of acid = 20 / 7413.1 = 0.0027 mili moles